# Towards a Method for LLM-enabled Indoor Navigation thanks: This work was founded by European Union - Next Generation EU, in the context of The National Recovery and Resilience Plan, Investment Partenariato Esteso PE8 ”Conseguenze e sfide dell’invecchiamento”, Project Age-IT, CUP: B83C22004880006

### v1

@@ Abstract
Indoor navigation presents unique challenges due to complex layouts, lack of GPS signals, and accessibility concerns. Existing solutions often struggle with real-time adaptability and user-specific needs. In this work, we explore the potential of a Large Language Model (LLM), i.e., ChatGPT, to generate natural, context-aware navigation instructions from indoor map images. We design and evaluate test cases across different real-world environments, analyzing the effectiveness of LLMs in interpreting spatial layouts, handling user constraints, and planning efficient routes. Our findings demonstrate the potential of LLMs for supporting personalized indoor navigation, with an average of 52% correct indications and a maximum of 62%. The results do not appear to depend on the complexity of the layout or the complexity of the expected path, but rather on the number of points of interest and the abundance of visual information, which negatively affect the performance.
@@ I Introduction
There is a growing need for a system capable of providing personalized, natural-language navigation assistance to address these limitations[4]. Such a system should interpret indoor maps and generate human-like, context-aware navigation instructions tailored to individual users. A system like this would also be highly beneficial for visually impaired individuals, as a detailed, natural-language description of the environment can enhance spatial awareness, helping them build a mental representation of their surroundings and navigate with greater confidence and independence. Integrating multimodal inputs, such as textual queries, voice commands, or sensor data, could provide more adaptive and accessible navigation experiences, ultimately fostering greater independence for individuals facing navigational barriers[5]. Beyond accessibility challenges, navigating unfamiliar indoor environments is another scenario where such a system would be beneficial. People visiting a new location, such as airports, shopping malls, hospitals, or office buildings, often struggle with orientation, requiring clear and accessible guidance to reach their destinations efficiently. A common approach to overcoming these challenges is to ask other people for directions, who typically provide landmark-based instructions (e.g., you will see a shoe shop, turn right, then you will encounter a phone charging area…) Given these challenges, leveraging Large Language Models (LLMs) for indoor navigation presents a promising alternative. LLMs can interpret complex multimodal inputs, reason over spatial relationships, and generate human-like natural language indications tailored to user-specific queries. This adaptability makes them well-suited for generating dynamic, human-like navigation instructions based on indoor maps[6]. In fact, recent advances in LLMs have opened new avenues for user-friendly navigation solutions for indoor and outdoor contexts, and preliminary studies have been performed leveraging LLMs to address this goal. For example, in [7], authors introduce TraveLLM, a system forpublic transit route planning. Their study compares multiple LLMs—including GPT-4, Claude 3, and Gemini—demonstrating that these models can integrate real-time data and user-specific constraints to form contextually aware navigation recommendations than conventional platforms (e.g., Google Maps). However, their study focuses on outdoor scenarios, leaving indoor navigation unexplored. Meanwhile, Zhang et al. [8] focus on improving navigation for people with visual impairments (PVI). NaviGPT integrates LiDAR-based obstacle detection, vibration feedback, and LLM-generated instructions to produce a real-time user experience. Unlike existing tools, e.g., Be My AI [^1] https://www.seeingai.com/ and Seeing AI  [^2] https://www.bemyeyes.com/blog/introducing-be-my-ai, which require multiple apps for obstacle detection, image recognition, or contextual directions, NaviGPT consolidates these capabilities into one cohesive system. This seamless integration offers continuous feedback without user distraction from frequent application switching. However, no evaluation of the system has been provided, and at the current stage, the proposal appears to mainly target outdoor environments. For indoor context, in [9], authors examine the development of geo-descriptions—textual narratives that describe the internal layout of a building. These authors combine classical GIS spatial analysis and LLM-based text generation to automate the creation of human-readable indoor maps, which is traditionally labour-intensive. However, the authors do not consider querying the system for navigation paths in the reconstructed textual representation of the building. Furthermore, the evaluation is performed on simplified cases.
@@ II Understanding user needs: Requirements for an inclusive experience
Indoor spaces vary widely in size, layout, and functionality, encompassing different types of destinations such as stores, restrooms, and emergency exits. Users often ask questions such as ”What can I find on this floor?” or ”Where is the nearest restroom?” highlighting the need for adaptable navigation solutions that can dynamically respond to different environments and user preferences. Among the many scenarios where such systems can be particularly beneficial is assisting users who may face additional navigation challenges, such as visual or hearing impairments, mobility constraints, or cognitive differences. While many existing systems are not designed to accommodate a broad spectrum of interaction needs, a more advanced approach could integrate personalized, multimodal assistance, ensuring that users receive navigation guidance in a format that best suits their preferences.
@@ III Experimental Set-up
For each map, five requests are made, each with a different category of difficulty. These were defined ad hoc based on a visual analysis of each map. The first three queries differ in terms of distance between starting and destination points, i.e., short (S), medium (M), and long (L). The other two queries, referred to as “specific requests” (Spec), explore other elements of complexity, which depend on each map. Specifically, we challenged the LLM with routes that have hidden or poorly marked IDs of points of interest, or with prompts in which the specific destination was not explicitly stated, e.g., ”Show me how to walk to the nearest dining area from gate number 26”. In this case, the query requires a more detailed search near gate 26 to locate the closest dining area. Table I reports examples of user prompts for the Bologna Airport.
@@
Figure 3: Correct path considered for ORIO-2 query (Table  I).
@@ IV-A Bergamo
The detailed results for Bergamo Airport are summarized in Table II. Over half of the indications are incorrect (60.78%), while 39.22%
IV-B Orio Center
@@
TABLE I: User prompt examples for the Bologna Airport.
@@
BOL-1 Medium Give me directions on how to walk to the Carrefour (15 on the map) from the Mondadori (7 on the map). Testing on a medium path
BOL-2 Short Give me directions on how to walk to the Vecchia Bologna Osteria (4 on the map) from the Carisbo (11 on the map). Testing on a short path.
@@
Table III presents the detailed results for Orio Center. Of 58 indications, 25 were incorrect (43.10%), and 33 were correct (56.90%). This indicates an improvement in terms of performance compared to the Bergamo case. The increased intricacy of Orio Center’s layout does not negatively impact the performance of the LLM. Again, the path length does not clearly correlate with performance, as the query of category L (ORIO-3) has the highest percentage of correct indications.
TABLE II: Results for Bergamo
@@
BER-1 (S) 7 (87.50%) 1 (12.50%)
BER-2 (L) 12 (54.55%) 10 (45.45%)
BER-3 (M) 5 (62.50%) 3 (37.50%)
BER-4 (Spec) 3 (60.00%) 2 (40.00%)
BER-5 (Spec) 4 (50.00%) 4 (50.00%)
Total 31 (60.78%) 20 (39.22%)
TABLE III: Results for Orio Center
Query Incorrect Correct
ORIO-1 (M) 3 (37.50%) 5 (62.50%)
ORIO-2 (M) 9 (69.23%) 4 (30.77%)
ORIO-3 (L) 3 (18.75%) 13 (81.25%)
ORIO-4 (S) 5 (71.43%) 2 (28.57%)
ORIO-5 (L) 5 (35.71%) 9 (64.29%)
Total 25 (43.10%) 33 (56.90%)
IV-C Bologna
Table IV presents the detailed results for Bologna. Out of 66 indications, 27 were incorrect (40.91%), and 39 were correct (59.09%). Here, it appears that the simpler layout, with fewer points of interest, leads to better accuracy. Again, no evident relation between the query category and the percentage of incorrect instructions can be observed.
TABLE IV: Results for Bologna
Query Incorrect Correct
BOL-1 (M) 4 (28.57%) 10 (71.43%)
BOL-2 (S) 8 (47.06%) 9 (52.94%)
BOL-3 (M) 6 (42.86%) 8 (57.14%)
BOL-4 (Spec) 5 (62.50%) 3 (37.50%)
BOL-5 (Spec) 4 (30.77%) 9 (69.23%)
Total 27 (40.91%) 39 (59.09%)
@@ IV-D Overall Performance
The overall performance across all locations is summarized in Table V. A total of 175 output indications were evaluated, with 83 incorrect (47.43%), and 92 correct (52.57%). This indicates that, while improvements are certainly required, the performance of the LLM is promising, also considering the complexity of the different scenarios and paths. While the best performance is obtained for the “simplest” map, i.e., Bologna Airport, challenging scenarios like Orio Center do not appear to affect the performance strongly. Instead, for a case like Bergamo, which also has a simple layout with a single corridor, performances are significantly lower. We argue that this is due to the high number of points of interest and more icons, which appear to overwhelm the LLMs. Our results suggest that it is not the layout or the complexity of the path that affects the performance, but rather the amount of information reported on the map. This can be improved by providing different, simplified views of a map to be used as input to the LLM to facilitate the task. Another phenomenon that we observed, which we can relate to the map’s visual appearance, is the types of errors made for the Orio Center case. In some situations, the LLM gave indications of routes that were not physically possible due to the presence of walls, which were not clearly recognizable due to the similar shades of colours on the map. We thus argue that an image with more contrasting colours could lead to better performance. This can be done in a pre-processing phase, together with removing the legend. It should be highlighted that the maps were all coloured, and in future works, we plan to evaluate the performance considering high-contrast greyscale images, which could simplify the task by discarding irrelevant colour information.
@@
Bergamo 31 (60.78%) 20 (39.22%)
Orio Center 25 (43.10%) 33 (56.90%)
@@ IV-E Comprehensive Evaluation of Incorrect Answer
In the comprehensive evaluation of the responses generated by the model during testing, it was discovered that 47.43% were incorrect. To understand the nature of these inaccuracies, a detailed analysis was undertaken. It became clear that these incorrect responses were not in accordance with the expected guidelines. However, this analysis also yielded valuable insights, revealing that the guidance provided does not negatively affect the user’s overall direction and progress despite the errors.
An example, taken from the response related to Figure  3, of insights categorized in this analysis as ’Partially Correct’ includes:
@@ V Conclusions
This paper evaluates the performance of ChatGPT in providing human-like indications to users in indoor environments represented through visual maps. We considered three different real-world map layouts with different characteristics, and five types of engineered user prompts of different complexity levels. Our preliminary results are promising, with an average of 47.43% correct indications, and a maximum of 52.57%. Furthermore, we noticed that the complexity of the user query and of the layout of the map do not appear to strongly influence the performance. Instead, these appear to be affected by the visual characteristics of the maps, e.g., the degree of contrast between colours and size and the information overload due to the density of points of interest. In future work, we plan to consolidate and improve these results with more systematic tests, including pre-processed maps, e.g., with increasing contrasts and simplified information. Furthermore, to go towards a more practical application of the envisioned approach, we plan to pre-process natural user queries so that these can be translated into engineered user prompts, similar to the ones used in our tests. These improvements aim to make the system more robust and user-friendly, paving the way for practical applications in real-world indoor navigation systems.

### v2

# Towards a Method for LLM-enabled Indoor Navigation (Arxiv) thanks: This work was founded by European Union - Next Generation EU, in the context of The National Recovery and Resilience Plan, Investment Partenariato Esteso PE8 ”Conseguenze e sfide dell’invecchiamento”, Project Age-IT, CUP: B83C22004880006

@@ Abstract
Indoor navigation presents unique challenges due to complex layouts, lack of GPS signals, and accessibility concerns. Existing solutions often struggle with real-time adaptability and user-specific needs. In this work, we explore the potential of a Large Language Model (LLM), i.e., ChatGPT, to generate natural, context-aware navigation instructions from indoor map images. We design and evaluate test cases across different real-world environments, analyzing the effectiveness of LLMs in interpreting spatial layouts, handling user constraints, and planning efficient routes. Our findings demonstrate the potential of LLMs for supporting personalized indoor navigation, with an average of 50.54% correct indications and a maximum of 77.78%. The results do not appear to depend on the complexity of the layout or the complexity of the expected path, but rather on the number of points of interest and the abundance of visual information, which negatively affect the performance.
@@ I Introduction
There is a growing need for a system capable of providing personalized, natural-language navigation assistance to address these limitations[4]. Such a system should interpret indoor maps and generate human-like, context-aware navigation instructions tailored to individual users. A system like this would also be highly beneficial for visually impaired individuals, as a detailed, natural-language description of the environment can enhance spatial awareness, helping them build a mental representation of their surroundings and navigate with greater confidence and independence. Integrating multimodal inputs, such as textual queries, voice commands, or sensor data, could provide more adaptive and accessible navigation experiences, ultimately fostering greater independence for individuals facing navigational barriers[5]. Beyond accessibility challenges, navigating unfamiliar indoor environments is another scenario where such a system would be beneficial. People visiting a new location, such as airports, shopping malls, hospitals, or office buildings, often struggle with orientation, requiring clear and accessible guidance to reach their destinations efficiently. A common approach to overcoming these challenges is to ask other people for directions, who typically provide landmark-based instructions (e.g., you will see a shoe shop, turn right, then you will encounter a phone charging area…). Given these challenges, leveraging Large Language Models (LLMs) for indoor navigation presents a promising alternative. LLMs can interpret complex multimodal inputs, reason over spatial relationships, and generate human-like natural language indications tailored to user-specific queries. This adaptability makes them well-suited for generating dynamic, human-like navigation instructions based on indoor maps[6]. In fact, recent advances in LLMs have opened new avenues for user-friendly navigation solutions for indoor and outdoor contexts, and preliminary studies have been performed leveraging LLMs to address this goal. For example, in [7], authors introduce TraveLLM, a system for public transit route planning. Their study compares multiple LLMs—including GPT-4, Claude 3, and Gemini—demonstrating that these models can integrate real-time data and user-specific constraints to form contextually aware navigation recommendations than conventional platforms (e.g., Google Maps). However, their study focuses on outdoor scenarios, leaving indoor navigation unexplored. Meanwhile, Zhang et al. [8] focus on improving navigation for people with visual impairments (PVI). NaviGPT integrates LiDAR-based obstacle detection, vibration feedback, and LLM-generated instructions to produce a real-time user experience. Unlike existing tools, e.g., Be My AI [^1] https://www.seeingai.com/ and Seeing AI  [^2] https://www.bemyeyes.com/blog/introducing-be-my-ai, which require multiple apps for obstacle detection, image recognition, or contextual directions, NaviGPT consolidates these capabilities into one cohesive system. This seamless integration offers continuous feedback without user distraction from frequent application switching. However, no evaluation of the system has been provided, and at the current stage, the proposal appears to mainly target outdoor environments. For indoor context, in [9], authors examine the development of geo-descriptions—textual narratives that describe the internal layout of a building. These authors combine classical GIS spatial analysis and LLM-based text generation to automate the creation of human-readable indoor maps, which is traditionally labour-intensive. However, the authors do not consider querying the system for navigation paths in the reconstructed textual representation of the building. Furthermore, the evaluation is performed on simplified cases.
@@ II Understanding user needs: Requirements for an inclusive experience
Indoor spaces vary widely in size, layout, and functionality, encompassing different types of destinations such as stores, restrooms, and emergency exits. Users often ask questions such as “What can I find on this floor?” or “Where is the nearest restroom?” highlighting the need for adaptable navigation solutions that can dynamically respond to different environments and user preferences. Among the many scenarios where such systems can be particularly beneficial is assisting users who may face additional navigation challenges, such as visual or hearing impairments, mobility constraints, or cognitive differences. While many existing systems are not designed to accommodate a broad spectrum of interaction needs, a more advanced approach could integrate personalized, multimodal assistance, ensuring that users receive navigation guidance in a format that best suits their preferences.
@@ III Experimental Set-up
For each map, five requests are made, each with a different category of difficulty. These were defined ad hoc based on a visual analysis of each map. The first three queries differ in terms of distance between starting and destination points, i.e., short (S), medium (M), and long (L). The other two queries, referred to as “specific requests” (Spec), explore other elements of complexity, which depend on each map. Specifically, we challenged the LLM with routes that have hidden or poorly marked IDs of points of interest, or with prompts in which the specific destination was not explicitly stated, e.g., “Show me how to walk to the nearest dining area from gate number 26”. In this case, the query requires a more detailed search near gate 26 to locate the closest dining area. Table III reports examples of user prompts for the Bologna Airport.
@@
Figure 3: Correct path considered for ORIO-2 query (Table  IV).
@@ IV-A Bergamo
The detailed results for Bergamo Airport are summarized in Table I. Over half of the indications are incorrect (62.26%), while 37.74% are correct.
IV-B Bologna
Table II presents the detailed results for Bologna. Out of 66 indications, 27 were incorrect (40.91%), and 39 were correct (59.09%). Here, it appears that the simpler layout, with fewer points of interest, leads to better accuracy. Again, no evident relation between the query category and the percentage of incorrect instructions can be observed.
TABLE I: Results for Bergamo
Query Incorrect Correct
BER-1 (S) 7 (87.50%) 1 (12.50%)
BER-2 (M) 7 (70.00%) 3 (30.00%)
BER-3 (L) 12 (54.55%) 10 (45.45%)
BER-4 (Spec) 3 (60.00%) 2 (40.00%)
BER-5 (Spec) 4 (50.00%) 4 (50.00%)
Total 33 (62.26%) 20 (37.74%)
TABLE II: Results for Bologna
Query Incorrect Correct
BOL-1 (S) 8 (47.06%) 9 (52.95%)
BOL-2 (M) 4 (28.57%) 10 (71.43%)
BOL-3 (L) 6 (42.86%) 8 (57.14%)
BOL-4 (Spec) 5 (62.50%) 3 (37.50%)
BOL-5 (Spec) 4 (30.77%) 9 (69.23%)
Total 27 (40.91%) 39 (59.09%)
IV-C Orio Center
@@
TABLE III: User prompt examples for the Bologna Airport.
@@
BOL-1 Short Give me directions on how to walk to the Vecchia Bologna Osteria (4 on the map) from the Carisbo (11 on the map) . Testing on a short path
BOL-2 Medium Give me directions on how to walk to the Carrefour (15 on the map) from the Mondadori (7 on the map). Testing on a medium path.
@@
Table IV presents the detailed results for Orio Center. Of 67 indications, 32 were incorrect (47.76%), and 35 were correct (52.24%). This indicates an improvement in terms of performance compared to the Bergamo case. The increased intricacy of Orio Center’s layout does not negatively impact the performance of the LLM. Again, the path length does not clearly correlate with performance, as the query of category L (ORIO-3) has the highest percentage of correct indications.
TABLE IV: Results for Orio Center
@@
ORIO-1 (S) 5 (71.43%) 2 (28.57%)
ORIO-2 (M) 5 (50.00%) 5 (50.00%)
ORIO-3 (L) 4 (22.22%) 14 (77.78%)
ORIO-4 (Spec) 13 (76.47%) 4 (23.53%)
ORIO-5 (Spec) 5 (33.33%) 10 (66.67%)
Total 32 (47.76%) 35 (52.24%)
@@ IV-D Overall Performance
The overall performance across all locations is summarized in Table V. A total of 186 output indications were evaluated, with 92 incorrect (49.64%), and 94 correct (50.54%). This indicates that, while improvements are certainly required, the performance of the LLM is promising, also considering the complexity of the different scenarios and paths. While the best performance is obtained for the “simplest” map, i.e., Bologna Airport, challenging scenarios like Orio Center do not appear to affect the performance strongly. Instead, for a case like Bergamo, which also has a simple layout with a single corridor, performances are significantly lower. We argue that this is due to the high number of points of interest and more icons, which appear to overwhelm the LLMs. Our results suggest that it is not the layout or the complexity of the path that affects the performance, but rather the amount of information reported on the map. This can be improved by providing different, simplified views of a map to be used as input to the LLM to facilitate the task. Another phenomenon that we observed, which we can relate to the map’s visual appearance, is the types of errors made for the Orio Center case. In some situations, the LLM gave indications of routes that were not physically possible due to the presence of walls, which were not clearly recognizable due to the similar shades of colours on the map. We thus argue that an image with more contrasting colours could lead to better performance. This can be done in a pre-processing phase, together with removing the legend. It should be highlighted that the maps were all coloured, and in future works, we plan to evaluate the performance considering high-contrast greyscale images, which could simplify the task by discarding irrelevant colour information.
@@
Bergamo 33 (62.26%) 20 (37.74%)
Orio Center 32 (47.76%) 35 (52.24%)
@@ IV-E Comprehensive Evaluation of Incorrect Answer
In the comprehensive evaluation of the responses generated by the model during testing, it was discovered that 49.46% were incorrect. To understand the nature of these inaccuracies, a detailed analysis was undertaken. It became clear that these incorrect responses were not in accordance with the expected guidelines. However, this analysis also yielded valuable insights, revealing that the guidance provided does not negatively affect the user’s overall direction and progress despite the errors.
An example, taken from the response related to Fig.  3, of insights categorized in this analysis as ’Partially Correct’ includes:
@@ V Conclusions
This paper evaluates the performance of ChatGPT in providing human-like indications to users in indoor environments represented through visual maps. We considered three different real-world map layouts with different characteristics, and five types of engineered user prompts of different complexity levels. Our preliminary results are promising, with an average of 50.54% correct indications, and a maximum of 77.78%. Furthermore, we noticed that the complexity of the user query and of the layout of the map do not appear to strongly influence the performance. Instead, these appear to be affected by the visual characteristics of the maps, e.g., the degree of contrast between colours and size and the information overload due to the density of points of interest. In future work, we plan to consolidate and improve these results with more systematic tests, including pre-processed maps, e.g., with increasing contrasts and simplified information. Furthermore, to go towards a more practical application of the envisioned approach, we plan to pre-process natural user queries so that these can be translated into engineered user prompts, similar to the ones used in our tests. These improvements aim to make the system more robust and user-friendly, paving the way for practical applications in real-world indoor navigation systems.